# The Landscape of Agentic Reinforcement Learning for LLMs: A Survey

### v1

@@ 3.7 Others
For long-horizon interactions, the central challenge is temporal credit assignment [109], where sparse and delayed feedback obscures the link between an agent’s actions and a distant outcome. Agentic RL directly confronts this by evolving both the learning signal and the optimization framework. One major approach is the (I) integration of process-based supervision with final outcome rewards. Rather than relying on a single reward at a trajectory’s conclusion, this paradigm uses auxiliary models or programmatic rules to evaluate the quality of intermediate steps, providing a denser and more immediate learning signal that guides the agent’s multi-turn strategy. For example, EPO [239], ThinkRM [240] and AgentPRM [241] introduce external reward models to provide step-wise signals for agents; in contrast, RLVMR [242] designs manually defined, programmatic rules to guide the intermediate supervision. A second, complementary strategy is to (II) extend preference optimization from single turns to multi-step segments. Techniques like Segment-level DPO (SDPO) [243] move beyond comparing isolated responses and instead construct preference data over entire conversational snippets or action sequences. This allows the model to directly learn how early decisions influence long-term success, thereby refining its ability to maintain strategic coherence in extended dialogues and complex tasks.
@@ 4 Agentic RL: The Task Perspective
Agentic RL manifests through a wide spectrum of concrete tasks that test and shape its evolving capabilities. This section surveys representative application domains where Agentic RL has demonstrated remarkable potential and unique challenges. We begin with search and information retrieval (Section 4.1), followed by code generation and software engineering (Section 4.2), and mathematical reasoning (Section 4.3). We then discuss its role in GUI navigation (Section 4.4), vision understanding tasks (Section 4.5) as well as VLM embodied interaction (Section 4.6). Beyond single-agent scenarios, we extend the perspective to multi-agent systems (Section 4.7) and conclude with other emerging domains (Section 4.8). Together, these applications highlight how agentic RL transitions from abstract paradigms into actionable, real-world problem solving, as illustrated in Figure 6.
Figure 6: The evolution tree of RL for domain-specific agents.
@@ 4.1 Search & Research Agent
Search has been central to extending LLMs with external knowledge, with Retrieval-Augmented Generation (RAG) as a widely used approach [244, 245]. The paradigm is now evolving beyond simple information retrieval towards creating autonomous agents capable of deep research: complex, multi-step processes that involve not just finding information but also performing in-depth analysis, synthesizing insights from numerous sources, and drafting comprehensive reports [104, 246]. This shift elevates the objective from answering queries to tackling complex research tasks. Early prompt-driven methods relied on brittle query strategies and manual engineering. While more recent works like Search-o1 [98] leverage large reasoning models for agentic, inference-time retrieval, and multi-agent systems such as DeepResearch [247] coordinate querying and summarization sub-agents, they still lack learning signals. These prompt-based methods lack any fine-tuning signal, leading to limited generalization and poor effectiveness in multi-turn settings that demand a tight loop of search, reasoning, and synthesis. These limitations have led to the adoption of reinforcement learning to directly optimize the end-to-end process of query generation and search–reasoning coordination for advanced research objectives. Table 4 presents the majority of works studied in this section. In the following, we will detail how RL empowers these agents.
@@
[…]
@@
[…]
OTC-PO [84] Internal / External Qwen2.5-3B/7B-Base/Instruct -
@@ 4.1.1 Open Source RL Methods
A major line of work builds on the RAG foundation but relies on real-time web search APIs as the external environment, using reinforcement learning to optimize query generation and multi-step reasoning. Early progress was spearheaded by DeepRetrieval [248], which framed one-shot query generation as a GRPO-trained policy and directly rewarded recall and relevance against live search results. Motivated by its gains, subsequent methods extended the paradigm into multi-turn, reasoning-integrated, and multi-modal search. Search-R1 [249] and DeepResearcher [253] integrates retrieved-token masking with outcome-based rewards to interleave query formulation and answer generation. R1-Searcher [250] employs a two-stage, cold-start PPO strategy—first learning when to invoke web search, then how to exploit it—while its successor R1-Searcher++ [251] adds supervised fine-tuning, internal-knowledge rewards to avoid redundancy, and dynamic memory for continual assimilation. ReSearch [99] pursues fully end-to-end PPO without supervised tool-use trajectories, while StepSearch [252] accelerates convergence on multi-hop QA by assigning intermediate step-level rewards. Atom-Searcher [257] is an agentic deep research framework that significantly improves LLM problem-solving by refining the reasoning process itself, not just the final outcome. WebDancer [97] leverages human browsing trajectory supervision plus RL fine-tuning to produce autonomous ReAct-style agents, excelling on GAIA [264] and WebWalkerQA [265]. WebThinker [254] embeds a Deep Web Explorer into a think-search-draft loop, aligning via DPO with human feedback to tackle complex report-generation. WebSailor [96] is a complete post-training methodology designed to teach LLM agents sophisticated reasoning for complex web navigation and information-seeking tasks. WebWatcher [255] further extends to multimodal search, combining visual-language reasoning, tool use, and RL to outperform text-only and multimodal baselines on BrowseComp-VL and VQA benchmarks. ASearcher [256] uses large-scale asynchronous reinforcement learning with synthesized QA data, enabling long-horizon search (40+ tool calls) and outperforming prior open-source methods. MiroMind Open Deep Research (MiroMind ODR) [258] is to build a high-performance, fully open-sourced, open-collaborative deep research ecosystem — with an agent framework, model, data, and training infra all fully accessible and open.
@@
However, these training methods that rely on external APIs face two major challenges: (1) the document quality of real-time internet document search is uncontrolled, and noisy information brings instability to the training process. (2) The API cost is too high and severely limits scalability. To enhance the efficiency, controllability and stability of training, some recent studies have used controllable simulated search engines such as LLM internal knowledge. For example, ZeroSearch [259] replaces live web retrieval with a pseudo search engine distilled from LLMs themselves, combining curriculum RL to gradually approach live-engine performance without issuing real queries. SSRL [260] takes this idea further: the agent performs entirely offline “self-search” during training, without explicit search engines, yet transfers seamlessly to online inference, where real APIs can still boost performance. Though still at an early stage, offline self-search enhances stability and scalability beyond API limits, pointing toward more self-reliant research agents.
@@ 4.1.2 Closed Source RL Methods
Despite progress in combining RAG and RL, most open source models still fail on OpenAI’s BrowseComp [266], a challenging benchmark that measures the ability of AI agents to locate hard-to-find information, revealing gaps in long-horizon planning, page-grounded tool use, and cross-source verification. In contrast, recent closed source systems are markedly stronger, having shifted from mere query optimization to fully autonomous research agents that navigate the open web, synthesize information from multiple sources, and draft comprehensive reports. This is likely due to the industry’s more powerful foundation models and the availability of more high-quality data. OpenAI Deep Research [103] achieves 51.5% pass@1 on BrowseComp. Other prototypes, Perplexity’s DeepResearch [246], Google Gemini’s DeepResearch [261], Kimi-Researcher [104], Grok AI DeepSearch [262], Doubao with Deep Think [263], combine RL-style fine-tuning with advanced tool integration and memory modules, ushering in a new era of interactive, iterative research assistants.
@@ 4.2 Code Agent
Code generation, or more broadly, software engineering, provides an ideal testbed for LLM-based agentic RL: execution semantics are explicit and verifiable, and automated signals (compilation, unit tests, and runtime traces) are readily available [267]. Early multi-agent frameworks (e.g., MetaGPT, AutoGPT, AgentVerse) coordinated roles through prompting without parameter updates, showcasing the promise of modular role allocation [268, 269, 270]. Initial RL for code, such as CodeRL, incorporated execution-based reward modeling and actor–critic training [271], catalyzing a wave of studies that exploit execution feedback to guide policy updates. Table 5 presents the majority of works studied in this section. In what follows, we structure the literature along increasing task complexity, progressing from single-turn code snippet generation (Section 4.2.1) to multi-turn codebase refinement (Section 4.2.2) and finally to automated software engineering (Section 4.2.3).
@@ 4.2.1 RL for Code Generation
Early research focused on relatively simple, single-round code generation (e.g., writing a function based on a prompt or solving a coding challenge in one go). RL was used to fine-tune LLMs to achieve higher levels of code correctness. This phase laid the foundation for subsequent large-scale software engineering.
@@
Methods in this class optimize directly for final correctness, typically measured by pass@k or unit-test success. Follow-ups extend outcome definitions to improve robustness: for example, AceCoder [272] rewards by execution pass rates to stabilize training on software maintenance–style prompts while staying within a single-turn setting. Beyond early actor–critic formulations, recent open-source efforts scale outcome-based RL on large pre-trained code models. DeepCoder-14B [273] fine-tunes a 14B model using distributed RL with unit-test pass rates as outcome rewards, achieving +8% Pass@1 gains (60.6% on LiveCodeBench) while matching much larger proprietary coders. A key to its stability is GRPO+, an improved proximal policy optimization variant that mitigates reward hacking and collapse by careful clipping and entropy modulation. RLTF employs an online RL loop that uses unit test results as multi-granularity reward signals, from coarse pass/fail outcomes to fine-grained fault localization, directly guiding code refinement [274]. CURE formalizes coder–tester co-evolution: a tester generates or evolves unit tests while a coder iteratively patches code; a reward-precision objective mitigates low-quality test effects during joint training [275]. Absolute Zero applies self-play RL without human data. It generates coding tasks for itself and uses execution outcomes as verifiable rewards to bootstrap reasoning ability [149]. Re:Form [276] leverages formal language-based reasoning with RL and automated verification to reduce human priors, enabling reliable program synthesis and surpassing strong baselines on formal verification tasks. In [277], authors propose a two-stage training pipeline—first fine-tuning for high code correctness, then applying an error-insensitive, efficiency-reward-driven online RL.
@@
To mitigate sparsity and credit assignment, process-aware designs provide intermediate supervision over compilation, partial execution, or reasoning steps. StepCoder [278] decomposes compilation and execution into step-level signals for shaping; Process Supervision-Guided Policy Optimization (PSGPO) [40] leverages intermediate error traces and process annotations for dense rewards; and CodeBoost [279] mines raw repositories to unify heterogeneous execution-derived signals, ranging from output correctness to error-message quality, under a single PPO framework. Further, PRLCoder [280] introduces process-supervised RL by constructing reward models that score each partial snippet: a teacher model mutates lines of reference solutions and assigns positive/negative signals based on compiler and test feedback. This fine-grained supervision yields faster convergence and +10.5% pass-rate improvements over the base model, illustrating how dense shaping at the line-level can guide code synthesis more effectively than outcome-only signals. o1-Coder [281] combines RL with Monte Carlo Tree Search, where the policy learns from exploration guided by test case rewards and gradually improves from pseudocode to executable code. Posterior-GRPO [282] rewards intermediate reasoning but gates credit by final test success to prevent speculative reward exploitation; Policy Filtration for RLHF [283] improves reward–correctness alignment by filtering low-confidence pairs before policy updates. Scaling preference supervision beyond costly human annotation has proven effective as well. CodeFavor [284] constructs CodePrefBench from code evolution histories, covering correctness, efficiency, security, and style to improve preference modeling and alignment. Focused-DPO [285] adapts preference-based RL by weighting preference optimization on error-prone regions of the code, making feedback more targeted and improving robustness across benchmarks.

### v4

@@ 3.7 Others
For long-horizon interactions, the central challenge is temporal credit assignment (Pignatelli et al., 2024), where sparse and delayed feedback obscures the link between an agent’s actions and a distant outcome. Agentic RL directly confronts this by evolving both the learning signal and the optimization framework. One major approach is the (I) integration of process-based supervision with final outcome rewards. Rather than relying on a single reward at a trajectory’s conclusion, this paradigm uses auxiliary models or programmatic rules to evaluate the quality of intermediate steps, providing a denser and more immediate learning signal that guides the agent’s multi-turn strategy. For example, EPO (Liu et al., 2025m), ThinkRM (Hong et al., 2025c), SPO (Guo et al., 2025c), and AgentPRM (Choudhury, 2025) introduce external reward models to provide step-wise signals for agents; in contrast, RLVMR (Zhang et al., 2025z) designs manually defined, programmatic rules to guide the intermediate supervision. A second, complementary strategy is to (II) extend preference optimization from single turns to multi-step segments. Techniques like Segment-level DPO (SDPO) (Kong et al., 2025) move beyond comparing isolated responses and instead construct preference data over entire conversational snippets or action sequences. This allows the model to directly learn how early decisions influence long-term success, thereby refining its ability to maintain strategic coherence in extended dialogues and complex tasks.
@@ 4 Agentic RL: The Task Perspective
Agentic RL manifests through a wide spectrum of concrete tasks that test and shape its evolving capabilities. This section surveys representative application domains where Agentic RL has demonstrated remarkable potential and unique challenges. We begin with search and information retrieval (Section 4.1), followed by code generation and software engineering (Section 4.2), and mathematical reasoning (Section 4.3). We then discuss its role in GUI navigation (Section 4.4), vision understanding tasks (Section 4.5), as well as VLM embodied interaction (Section 4.6). Beyond single-agent scenarios, we extend the perspective to multi-agent systems (Section 4.7) and conclude with other emerging domains (Section 4.8). Together, these applications highlight how Agentic RL transitions from abstract paradigms into actionable, real-world problem-solving, as illustrated in Figure 6.
Figure 6: The evolution tree of RL for domain-specific agents, illustrating the chronological progression of representative domains and methods.
@@ 4.1 Search & Research Agent
Search has been central to extending LLMs with external knowledge, with Retrieval-Augmented Generation (RAG) as a widely used approach (Gao et al., 2024; Fan et al., 2024). The paradigm is now evolving beyond simple information retrieval towards creating autonomous agents capable of deep research: complex, multi-step processes that involve not just finding information but also performing in-depth analysis, synthesizing insights from numerous sources, and drafting comprehensive reports (Kimi, 2025; Perplexity, 2025). This shift elevates the objective from answering queries to tackling complex research tasks. Early prompt-driven methods relied on brittle query strategies and manual engineering. While more recent works like Search-o1 (Li et al., 2025i) leverage large reasoning models for agentic, inference-time retrieval, and multi-agent systems such as DeepResearch (Zhang et al., 2025r) coordinate querying and summarization sub-agents, they still lack learning signals. These prompt-based methods lack any fine-tuning signal, leading to limited generalization and poor effectiveness in multi-turn settings that demand a tight loop of search, reasoning, and synthesis. These limitations have led to the adoption of reinforcement learning to directly optimize the end-to-end process of query generation and search–reasoning coordination for advanced research objectives. Table 4 presents the majority of works studied in this section. In the following, we will detail how RL empowers these agents.
@@
DeepRetrieval (Jiang et al., 2025c) External Qwen2.5-3B-Instruct, Llama-3.2-3B-Instruct GitHub
Search-R1 (Jin et al., 2025a) External Qwen2.5-3B/7B-Base/Instruct GitHub
R1-Searcher (Song et al., 2025a) External Qwen2.5-7B, Llama3.1-8B-Instruct GitHub
R1-Searcher++ (Song et al., 2025b) External Qwen2.5-7B-Instruct GitHub
ReSearch (Chen et al., 2025d) External Qwen2.5-7B/32B-Instruct GitHub
StepSearch (Wang et al., 2025w) External Qwen2.5-3B/7B-Base/Instruct GitHub
DeepResearcher (Zheng et al., 2025e) External Qwen2.5-7B-Instruct GitHub
WebDancer (Wu et al., 2025a) External Qwen2.5-7B/32B, QWQ-32B GitHub
WebThinker (Li et al., 2025j) External QwQ-32B, DeepSeek-R1-Distilled-Qwen, Qwen2.5-32B GitHub
WebSailor (Li et al., 2025f) External Qwen2.5-3B/7B/32B/72B GitHub
WebWatcher (Geng et al., 2025) External Qwen2.5-VL-7B/32B GitHub
WebShaper (Tao et al., 2025) External Qwen-2.5-32B/72B, QwQ-32B GitHub
ASearcher (Gao et al., 2025c) External Qwen2.5-7B/14B, QwQ-32B GitHub
Atom-Searcher (Deng et al., 2025b) External Qwen2.5-7B-Instruct GitHub
MiroMind Open Deep Research (MiroMind Team, 2025) External - Website
SimpleDeepResearcher (Sun et al., 2025b) External QwQ-32B GitHub
AWorld (Yu et al., 2025a) External Qwen3-32B GitHub
SFR-DeepResearch (Nguyen et al., 2025b) External QwQ-32B, Qwen3-8B, GPT-oss-20b -
ZeroSearch (Sun et al., 2025a) Internal Qwen2.5-3B/7B-Base/Instruct GitHub
SSRL (Fan et al., 2025b) Internal Qwen2.5, Llama-3.2/Llama-3.1, Qwen3 GitHub
@@
OpenAI Deep Research (OpenAI, 2025) External OpenAI Models Website
Perplexity’s DeepResearch (Perplexity, 2025) External - Website
Google Gemini’s DeepResearch (Google, 2025) External Gemini Website
Kimi-Researcher (Kimi, 2025) External Kimi K2 Website
Grok AI DeepSearch (x.ai, 2025) External Grok3 Website
Doubao with Deep Think (Doubao, 2025) External Doubao Website
Manus WideResearch External - Website
@@ 4.1.1 Open Source RL Methods
A major line of work builds on the RAG foundation but relies on real-time web search APIs as the external environment, using reinforcement learning to optimize query generation and multi-step reasoning. Early progress was spearheaded by DeepRetrieval (Jiang et al., 2025c), which framed one-shot query generation as a GRPO-trained policy and directly rewarded recall and relevance against live search results. Motivated by its gains, subsequent methods extended the paradigm into multi-turn, reasoning-integrated, and multi-modal search. Search-R1 (Jin et al., 2025a) and DeepResearcher (Zheng et al., 2025e) integrate retrieved-token masking with outcome-based rewards to interleave query formulation and answer generation. AutoRefine (Shi et al., 2025b) further advances this trajectory by inserting refinement phases between successive search calls, using GRPO to reward not only answer correctness but also retrieval quality, enabling agents to iteratively filter and structure noisy evidence during long-horizon reasoning. R1-Searcher (Song et al., 2025a) employs a two-stage, cold-start PPO strategy—first learning when to invoke web search, then how to exploit it—while its successor R1-Searcher++ (Song et al., 2025b) adds supervised fine-tuning, internal-knowledge rewards to avoid redundancy, and dynamic memory for continual assimilation. ReSearch (Chen et al., 2025d) pursues fully end-to-end PPO without supervised tool-use trajectories, while StepSearch (Wang et al., 2025w) accelerates convergence on multi-hop QA by assigning intermediate step-level rewards. Atom-Searcher (Deng et al., 2025b) is an agentic deep research framework that significantly improves LLM problem-solving by refining the reasoning process itself, not just the final outcome. WebDancer (Wu et al., 2025a) leverages human browsing trajectory supervision plus RL fine-tuning to produce autonomous ReAct-style agents, excelling on GAIA (Mialon et al., 2024) and WebWalkerQA (Wu et al., 2025b). WebThinker (Li et al., 2025j) embeds a Deep Web Explorer into a think-search-draft loop, aligning via DPO with human feedback to tackle complex report-generation. WebSailor (Li et al., 2025f) is a complete post-training methodology designed to teach LLM agents sophisticated reasoning for complex web navigation and information-seeking tasks. WebWatcher (Geng et al., 2025) further extends to multimodal search, combining visual-language reasoning, tool use, and RL to outperform text-only and multimodal baselines on BrowseComp-VL and VQA benchmarks. ASearcher (Gao et al., 2025c) uses large-scale asynchronous reinforcement learning with synthesized QA data, enabling long-horizon search (40+ tool calls) and outperforming prior open-source methods. MiroMind Open Deep Research (MiroMind ODR) (MiroMind Team, 2025) aims to build a high-performance, fully open-sourced, open-collaborative deep research ecosystem — with an agent framework, model, data, and training infrastructure all fully accessible and open.
@@
However, these training methods that rely on external APIs face two major challenges: (1) the document quality of real-time internet document searching is uncontrolled, and noisy information brings instability to the training process. (2) The API cost is too high and severely limits scalability. To enhance the efficiency, controllability, and stability of training, some recent studies have used controllable simulated search engines such as LLM internal knowledge. For example, ZeroSearch (Sun et al., 2025a) replaces live web retrieval with a pseudo search engine distilled from LLMs themselves, combining curriculum RL to gradually approach live-engine performance without issuing real queries. SSRL (Fan et al., 2025b) takes this idea further: the agent performs entirely offline “self-search” during training, without explicit search engines, yet transfers seamlessly to online inference, where real APIs can still boost performance. Though still at an early stage, offline self-search enhances stability and scalability beyond API limits, pointing toward more self-reliant research agents.
@@ 4.1.2 Closed Source RL Methods
Industrial Research Agents.
Despite progress in combining RAG and RL, most open-source models still fail on OpenAI’s BrowseComp (Wei et al., 2025b), a challenging benchmark that measures the ability of AI agents to locate hard-to-find information, revealing gaps in long-horizon planning, page-grounded tool use, and cross-source verification. In contrast, recent closed source systems are markedly stronger, having shifted from mere query optimization to fully autonomous research agents that navigate the open web, synthesize information from multiple sources, and draft comprehensive reports. This is likely due to the industry’s more powerful foundation models and the availability of more high-quality data. OpenAI Deep Research (OpenAI, 2025) achieves 51.5% pass@1 on BrowseComp. Other prototypes, Perplexity’s DeepResearch (Perplexity, 2025), Google Gemini’s DeepResearch (Google, 2025), Kimi-Researcher (Kimi, 2025), Grok AI DeepSearch (x.ai, 2025), Doubao with Deep Think (Doubao, 2025), combine RL-style fine-tuning with advanced tool integration and memory modules, ushering in a new era of interactive, iterative research assistants.
Case Study: OpenAI Deep Research.
Deep Research provides a concrete example of how capabilities from Section 3 combine with the RL-shaped search strategies. The agent begins with long-horizon multi-step reasoning and planning, decomposing a user request into sub-goals. It then performs RL-shaped web search: issuing queries, selecting which pages to open, and refining its search trajectory. These search policies are shaped during training using research-oriented benchmarks such as BrowseComp (Wei et al., 2025b). Throughout the process, the agent maintains persistent memory in the form of scratchpad notes and performs cross-source verification before synthesis. These capabilities—reasoning, planning, tool use, memory, and verification—are coupled with RL-shaped control decisions over search depth, branch selection, and evidence integration, forming a unified research agent.
@@ 4.2 Code Agent
Code generation, or more broadly, software engineering, provides an ideal testbed for LLM-based Agentic RL: execution semantics are explicit and verifiable, and automated signals (compilation, unit tests, and runtime traces) are readily available (Dong et al., 2025b). Early multi-agent frameworks (e.g., MetaGPT, AutoGPT, AgentVerse) coordinated roles through prompting without parameter updates, showcasing the promise of modular role allocation (Hong et al., 2024b; Gravitas, 2023; Chen et al., 2024e). Initial RL for code, such as CodeRL, incorporated execution-based reward modeling and actor–critic training (Le et al., 2022), catalyzing a wave of studies that exploit execution feedback to guide policy updates. Table 5 presents the majority of works studied in this section. We structure the literature along increasing task complexity, progressing from code generation (Section 4.2.1) to code refinement (Section 4.2.2) and software engineering (Section 4.2.3).
@@ 4.2.1 RL for Code Generation
Early research focused on relatively simple, single-round code generation (e.g., completing a function or solving a coding challenge in one go), which lays the foundation for subsequent large-scale software engineering.
@@
Methods in this class optimize directly for final correctness, typically measured by pass@k or unit-test success. AceCoder (Zeng et al., 2025b) introduces a data-efficient RLHF pipeline for code generation, constructing large-scale preference pairs from existing code fragments to train a reward model via Bradley–Terry loss, which then guides RFT on the synthesized dataset. Beyond early actor–critic formulations, recent open-source efforts scale outcome-based RL on large pre-trained code models. DeepCoder-14B (Luo et al., 2025c) stabilizes GRPO training via iterative context lengthening and DAPO-inspired filtering, and employs a sparse Outcome Reward Model (ORM) to prevent reward hacking on curated coding data. RLTF employs an online RL loop that uses unit test results as multi-granularity reward signals, from coarse pass/fail outcomes to fine-grained fault localization, directly guiding code refinement (Liu et al., 2023b). CURE formalizes coder–tester co-evolution: a tester generates or evolves unit tests while a coder iteratively patches code; a reward-precision objective mitigates low-quality test effects during joint training (Wang et al., 2025q). Absolute Zero applies self-play RL without human data. It generates coding tasks for itself and uses execution outcomes as verifiable rewards to bootstrap reasoning ability (Zhao et al., 2025a). Re:Form (Yan et al., 2025a) leverages formal language-based reasoning with RL and automated verification to reduce human priors, enabling reliable program synthesis and surpassing strong baselines on formal verification tasks. In (Feng et al., 2025c), the authors propose a two-stage training pipeline: first fine-tuning for a high-correctness baseline, then performing efficiency-driven online RL optimization.
@@
To mitigate sparsity and credit assignment, several works design process-level supervision by integrating compilation and execution feedback. StepCoder (Dou et al., 2024) decomposes compilation and execution into step-level signals for shaping; Process Supervision-Guided Policy Optimization (PSGPO) (Dai et al., 2025) leverages intermediate error traces and process annotations for dense rewards; and CodeBoost (Wang et al., 2025n) mines raw repositories to unify heterogeneous execution-derived signals, ranging from output correctness to error-message quality, under a single PPO framework. Further, PRLCoder (Ye et al., 2025b) introduces process-supervised RL by constructing reward models that score each partial snippet: a teacher model mutates lines of reference solutions and assigns positive/negative signals based on compiler and test feedback. This fine-grained supervision yields faster convergence and +10.5% pass-rate improvements over the base model, illustrating how dense shaping at the line-level can guide code synthesis more effectively than outcome-only signals. o1-Coder (Zhang et al., 2024d) combines RL with Monte Carlo Tree Search, where the policy learns from exploration guided by test case rewards and gradually improves from pseudocode to executable code. Posterior-GRPO (Fan et al., 2025a) rewards intermediate reasoning but gates credit by final test success to prevent speculative reward exploitation; Policy Filtration for RLHF (Zhang et al., 2025c) improves reward-correctness alignment by filtering low-confidence pairs before policy updates. Scaling preference supervision beyond costly human annotation has proven effective as well. CodeFavor (Liu et al., 2024a) constructs CodePrefBench from code evolution histories, covering correctness, efficiency, security, and style to improve preference modeling and alignment. Focused-DPO (Zhang et al., 2025o) adapts preference-based RL by weighting preference optimization on error-prone regions of the code, making feedback more targeted and improving robustness across benchmarks. Yang et al. (2025f) studies how RL-trained small-scale agents surpass large-scale prompt-based models in MLE environments via duration-aware gradient updates in a distributed asynchronous RL.